Assume you have a lawn that is 45 feet by 25 feet. Further, assume you buy a trampoline (a circular one) that is 15 feet in diameter. How much of the lawn is NOT under the tramp?
1
Expert's answer
2011-07-01T10:41:21-0400
The area of the trampoline is
π d2/4 = 255/4π = 63.75 π = 200.28 ft2
The area of the lawn
45 * 25 = 1125 ft2.
1125-200.28 = 924.72 ft2 of the lawn is not under the trump.